OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 74 and the last term is –100? a.–468 b.–442 c. –416 d. –390

OpenStudy (amistre64):

the general formula is:\[\frac{n(first+last)}{2}\]

OpenStudy (anonymous):

n(74-100)/2 -26n/2 do i go further & simply ?

OpenStudy (amistre64):

so far so good; but yes go further. n=30 in this case some other ideas i have are 30/2 = 15 74-100 = 26 since the last digit of 6*5 = 30 is zero; it should end in a 0

OpenStudy (anonymous):

so it's -390

OpenStudy (amistre64):

i didnt do any thorough mathing, but all signs point to yes :)

OpenStudy (anonymous):

Thanks for the help